# Disentangling Reasoning and Knowledge in Medical Large Language Models

### v1

@@ 5 Discussion and Conclusion
In this work, we disentangled reasoning from knowledge across 11 biomedical QA benchmarks and found that only 32% of questions require complex reasoning, challenging the widespread assumption that these benchmarks robustly assess clinical reasoning. These datasets are frequently cited as reasoning evaluations, and recent biomedical models often highlight improved performance on them as evidence of enhanced reasoning ability. However, our stratified analysis suggests that such gains are more likely attributable to improved factual recall than to genuine reasoning. We observed consistent performance gaps across biomedical LLMs between reasoning- and knowledge-heavy questions. Robustness testing with adversarial traces further exposed the limited self-correction capabilities of biomedical models, while general-domain models such as R1 and o4-mini demonstrated greater resilience. These findings highlight a critical shortcoming: despite instruction tuning and reasoning supervision, current biomedical models remain brittle when confronted with misleading or uncertain starting points. This brittleness underscores the absence of robust reasoning behavior—deliberate, analytical re-evaluation of flawed hypotheses—which is essential for safe and reliable clinical deployment. While our results offer early evidence that RL on high-reasoning examples can enhance both reasoning and knowledge accuracy—as well as adversarial robustness—substantial room for improvement remains. Incorporating additional reasoning-rich supervision, such as clinical case reports, and designing training setups that simulate adversarial traces and promote iterative self-correction may further strengthen reasoning robustness.
@@
While our work provides new insights into evaluating and improving medical reasoning in LLMs, several limitations remain. First, although we develop an automated classifier to distinguish reasoning from knowledge-heavy questions, cleanly disentangling these categories is inherently difficult—even expert annotators often disagree. Nonetheless, we show that even an approximate separation reveals consistent performance gaps that meaningfully inform model capabilities. Second, our evaluation focuses on three recent biomedical reasoning models. A broader study incorporating additional biomedical models, reasoning datasets, and distillation strategies from general-domain LLMs could offer a more comprehensive perspective. Similarly, while we present preliminary findings on reinforcement learning, a more systematic comparison between SFT and RL-based approaches remains an important direction for future work. Finally, our robustness testing experiments highlight the vulnerability of biomedical models to adversarial prompting, but we do not explore mitigation strategies. Developing robust self-correction and error recovery mechanisms remains an open challenge, and we leave this important direction for future work.
@@
Appendix A Additional Training Details
@@
FSDP Full shard, no CPU offload, rank0 init
@@ Appendix B Additional Data Curation Details
Example of Adversarial Trace Prefill for Llama-3.1-8B-Instruct Inference <|eot_id|><|start_header_id|>user<|end_header_id|> You are a biomedical reasoning model. You must think step-by-step and reason carefully about the following question before answering. You must select the correct option from the given options (A, B, C, D, etc.). Your response must conclude with the correct option in the format: <answer> A </answer>. What is the impact of knocking out LIG1 on CTG somatic instability in the context of Myotonic dystrophy? A. Increased instability B. Contraction of repeat length C. Reduced instability D. No impact E. Knockout is lethal<|eot_id|><|start_header_id|>assistant<|end_header_id|> Perhaps A. Increased instability is correct, but I should examine this more carefully before deciding.
Figure S8: Example of pre-filling model input with an adversarial trace to simulate incorrect reasoning. Used to evaluate the model’s ability to recover and self-correct.
@@
Dataset Knowledge-heavy Reasoning-heavy Total
HLE (Biomed) 115 (78.2%) 32 (21.8%) 147
MedBullets (4 options) 7 (2.3%) 301 (97.7%) 308
MedBullets (5 options) 7 (2.3%) 301 (97.7%) 308
MedMCQA 3826 (91.5%) 357 (8.5%) 4183
MedQA-USMLE 440 (34.6%) 833 (65.4%) 1273
MedXpertQA 777 (31.7%) 1673 (68.3%) 2450
MMLU (Health/Biology) 1356 (88.3%) 179 (11.7%) 1535
PubMedQA 958 (95.8%) 42 (4.2%) 1000
GPQA (Medical) 375 (96.2%) 15 (3.8%) 390
Lancet 281 (68.2%) 131 (31.8%) 412
NEJM 329 (54.6%) 274 (45.4%) 603
Total 8471 (67.2%) 4138 (32.8%) 12609
Model Knowledge Reasoning Overall
Qwen w/ self-consistency (64) 53.051.9 - ${}_{\text{51.9 - 54.0}}$ 35.734.3 - ${}_{\text{34.3 - 37.1}}$ 47.346.4 - ${}_{\text{46.4 - 48.2}}$
Qwen w/ CoT 54.553.5 - ${}_{\text{53.5 - 55.6}}$ 36.034.6 - ${}_{\text{34.6 - 37.5}}$ 48.447.6 - ${}_{\text{47.6 - 49.3}}$
Qwen w/ retrieval on medical o1 54.953.9 - ${}_{\text{53.9 - 56.0}}$ 37.936.5 - ${}_{\text{36.5 - 39.3}}$ 49.348.4 - ${}_{\text{48.4 - 50.1}}$
Qwen w/ medical o1 SFT w/o Reasoning 55.354.3 - ${}_{\text{54.3 - 56.4}}$ 39.237.8 - ${}_{\text{37.8 - 40.5}}$ 50.049.1 - ${}_{\text{49.1 - 50.9}}$
Qwen w/ medical o1 SFT 54.653.5 - ${}_{\text{53.5 - 55.6}}$ 38.937.4 - ${}_{\text{37.4 - 40.5}}$ 49.548.6 - ${}_{\text{48.6 - 50.3}}$
Qwen w/ retrieval on MedReason 54.753.6 - ${}_{\text{53.6 - 55.7}}$ 40.739.1 - ${}_{\text{39.1 - 42.0}}$ 50.149.2 - ${}_{\text{49.2 - 50.9}}$
Qwen w/ MedReason SFT w/o Reasoning 55.954.9 - ${}_{\text{54.9 - 56.8}}$ 40.338.9 - ${}_{\text{38.9 - 41.9}}$ 50.849.9 - ${}_{\text{49.9 - 51.6}}$
Qwen w/ MedReason SFT 54.453.3 - ${}_{\text{53.3 - 55.4}}$ 38.336.9 - ${}_{\text{36.9 - 39.7}}$ 49.148.2 - ${}_{\text{48.2 - 49.9}}$
Qwen w/ retrieval on m23k 55.754.6 - ${}_{\text{54.6 - 56.8}}$ 38.537.0 - ${}_{\text{37.0 - 40.0}}$ 50.149.2 - ${}_{\text{49.2 - 50.9}}$
Qwen w/ m23k SFT w/o Reasoning 52.051.0 - ${}_{\text{51.0 - 53.0}}$ 36.835.5 - ${}_{\text{35.5 - 38.3}}$ 47.046.1 - ${}_{\text{46.1 - 47.9}}$
Qwen w/ m23k SFT 55.954.8 - ${}_{\text{54.8 - 57.0}}$ 41.339.8 - ${}_{\text{39.8 - 42.8}}$ 51.150.3 - ${}_{\text{50.3 - 52.0}}$
Table S2: Accuracy across knowledge, reasoning, and overall categories for Qwen2.5-7B-Instruct variants. Confidence intervals (CI) are shown as subscript ranges. “SFT Baselines (Final Answer Only)” refers to models trained without reasoning traces. All variants show a consistent performance gap between reasoning and knowledge, underscoring the difficulty of reasoning-heavy questions. Models trained on the m23k dataset perform best, especially when reasoning traces are included.
@@
DeepSeek-R1-Distill-Llama-8B 55.9 49.7 74.8 58.6 57.1 57.1 16.0 6.1 60.8 53.0 46.5
w/ self-consistency (64) 53.2 51.0 77.0 54.3 42.9 42.9 13.3 13.9 45.3 52.3 56.2
w/ CoT 70.7 58.9 77.2 60.6 57.1 42.9 16.1 8.7 52.0 60.5 57.8
w/ retrieval on medical o1 61.8 57.8 79.0 60.4 71.4 57.1 16.7 10.4 53.6 58.0 55.0
w/ retrieval on MedReason 68.9 59.3 79.3 60.4 42.9 57.1 20.1 12.2 44.8 58.7 59.6
w/ retrieval on m23k 65.2 59.5 75.9 59.1 85.7 42.9 17.9 13.9 50.1 59.8 59.0
w/ medical o1 SFT Baseline 60.2 58.1 76.4 58.6 57.1 71.4 16.1 13.0 38.7 59.8 59.3
w/ MedReason SFT Baseline 61.8 58.9 67.3 58.1 42.9 42.9 17.2 19.1 48.0 60.5 59.3
w/ m23k SFT Baseline 50.2 54.5 76.3 55.1 71.4 57.1 17.0 19.1 42.7 57.3 53.2
w/ medical o1 SFT 69.3 58.5 76.2 60.8 57.1 57.1 15.1 10.4 52.8 58.0 60.5
w/ medical R1-Distill SFT 78.0 63.1 76.1 63.9 57.1 85.7 18.9 12.2 48.8 63.7 61.7
w/ MedReason SFT 65.9 56.9 76.0 57.9 57.1 71.4 15.2 15.7 44.5 56.9 59.3
w/ m23k SFT 78.6 63.4 77.0 64.0 71.4 57.1 19.4 12.2 47.2 64.8 63.8
                          Reasoning Accuracy
DeepSeek-R1-Distill-Llama-8B 48.1 46.2 76.2 41.9 44.5 38.9 14.1 9.4 33.3 58.8 46.7
w/ self-consistency (64) 53.4 57.1 83.3 55.3 47.8 40.9 13.7 3.1 13.3 58.8 55.8
w/ CoT 60.7 60.2 83.3 65.4 56.5 50.2 13.8 6.2 26.7 58.0 65.3
w/ retrieval on medical o1 59.1 58.0 85.7 55.9 55.5 49.8 16.4 9.4 46.7 61.1 60.6
w/ retrieval on MedReason 62.4 59.4 88.1 61.5 57.1 55.8 20.0 3.1 26.7 60.3 59.9
w/ retrieval on m23k 61.9 57.1 88.1 62.0 53.5 50.2 16.3 15.6 60.0 59.5 63.1
w/ medical o1 SFT Baseline 61.0 61.6 83.3 59.8 54.8 47.2 17.6 6.2 66.7 60.3 59.5
w/ MedReason SFT Baseline 61.0 60.5 78.6 58.1 61.8 54.8 22.1 12.5 33.3 60.3 61.7
w/ m23k SFT Baseline 49.3 54.9 81.0 53.6 52.5 46.8 19.4 6.2 66.7 55.7 51.1
w/ medical o1 SFT 61.2 63.9 90.5 62.6 56.8 51.8 16.1 6.2 60.0 64.9 59.5
w/ medical R1-Distill SFT 72.1 65.5 88.1 75.4 66.8 58.5 17.5 12.5 46.7 61.8 68.2
w/ MedReason SFT 62.3 54.6 92.9 63.1 55.5 52.8 15.7 3.1 6.7 59.5 61.7
w/ m23k SFT 71.1 60.2 88.1 71.5 65.1 59.5 19.5 6.2 26.7 66.4 67.2
                          Overall Accuracy
DeepSeek-R1-Distill-Llama-8B 50.8 49.4 74.9 56.7 44.8 39.3 14.7 6.8 59.7 54.9 46.6
w/ self-consistency (64) 53.3 51.5 77.3 54.4 47.7 40.9 13.6 11.6 44.1 54.4 56.1
w/ CoT 64.2 59.0 77.5 61.2 56.5 50.0 14.5 8.2 51.0 59.7 61.2
w/ retrieval on medical o1 60.0 57.8 79.3 59.9 55.8 50.0 16.5 10.2 53.3 59.0 57.5
w/ retrieval on MedReason 64.7 59.3 79.7 60.5 56.8 55.8 20.0 10.2 44.1 59.2 59.7
w/ retrieval on m23k 63.1 59.3 76.4 59.4 54.2 50.0 16.8 14.3 50.5 59.7 60.9
w/ medical o1 SFT Baseline 60.7 58.4 76.7 58.7 54.9 47.7 17.1 11.6 39.7 60.0 59.4
w/ MedReason SFT Baseline 61.3 59.0 67.8 58.1 61.4 54.5 20.6 17.7 47.4 60.4 60.4
w/ m23k SFT Baseline 49.6 54.6 76.5 54.9 52.9 47.1 18.6 16.3 43.6 56.8 52.2
w/ medical o1 SFT 64.0 59.0 76.8 61.0 56.8 51.9 15.8 9.5 53.1 60.2 60.0
w/ medical R1-Distill SFT 74.2 63.4 76.6 65.2 66.6 59.1 17.9 12.2 48.7 63.1 64.7
w/ MedReason SFT 63.6 56.7 76.7 58.5 55.5 53.2 15.6 12.9 43.1 57.8 60.4
w/ m23k SFT 73.7 63.2 77.5 64.9 65.3 59.4 19.5 10.9 46.4 65.3 65.3
Model MQA MCQA PQA MMLU MB4 MD5 MDX HLE GPQA Lct. NEJM
                          Knowledge Accuracy
DeepSeek-R1-Distill-Qwen-7B 37.7 38.9 74.1 49.9 57.1 28.6 12.4 11.3 61.3 33.8 31.6
Qwen w/ self-consistency (64) 51.4 55.3 73.5 57.4 57.1 57.1 12.9 20.0 50.9 58.0 54.4
w/ CoT 58.9 55.8 75.5 62.3 57.1 57.1 12.6 17.4 52.5 58.0 51.4
w/ retrieval on medical o1 57.5 56.0 76.5 64.5 71.4 71.4 16.1 13.0 45.6 53.4 52.6
w/ retrieval on MedReason 58.0 55.9 76.3 62.4 42.9 57.1 17.8 16.5 47.7 53.7 50.2
w/ retrieval on m23k 63.2 56.6 77.2 62.7 57.1 57.1 17.0 13.0 48.0 57.3 57.1
w/ medical o1 SFT Baseline 55.7 56.5 76.9 63.7 71.4 57.1 14.9 18.3 44.3 59.8 60.2
w/ MedReason SFT Baseline 57.7 57.5 77.7 63.3 57.1 71.4 13.6 19.1 48.8 59.4 57.8
w/ m23k SFT Baseline 45.2 52.4 76.3 57.7 42.9 71.4 18.8 18.3 51.7 53.7 52.0
w/ medical o1 SFT 63.0 55.3 73.1 62.2 42.9 71.4 13.4 12.2 54.7 58.4 58.7
w/ medical R1-Distill SFT 69.1 58.7 75.5 64.9 57.1 71.4 15.3 15.7 46.9 58.0 58.4
w/ MedReason SFT 62.5 56.0 76.5 61.5 71.4 42.9 13.1 14.8 44.5 54.4 52.6
w/ m23k SFT 67.5 58.9 73.0 62.0 57.1 57.1 15.6 14.8 42.9 57.3 53.5
                          Reasoning Accuracy
DeepSeek-R1-Distill-Qwen-7B 27.6 36.4 83.3 23.5 30.6 25.9 11.7 9.4 40.0 35.1 31.0
w/ self-consistency (64) 52.2 52.9 85.7 47.5 48.8 39.9 12.4 12.5 60.0 58.8 61.7
w/ CoT 54.0 50.4 88.1 58.1 48.8 40.9 11.8 9.4 33.3 61.8 58.8
w/ retrieval on medical o1 54.1 51.5 83.3 60.9 51.5 39.9 15.9 12.5 40.0 62.6 56.9
w/ retrieval on MedReason 57.3 54.1 92.9 58.1 58.8 48.8 18.6 3.1 40.0 61.1 55.1
w/ retrieval on m23k 54.4 54.9 90.5 62.0 55.5 42.9 15.3 9.4 33.3 65.6 54.7
w/ medical o1 SFT Baseline 55.1 54.1 85.7 57.0 51.5 46.5 15.2 9.4 66.7 71.8 64.2
w/ MedReason SFT Baseline 59.2 53.2 85.7 58.7 55.5 50.8 15.2 15.6 66.7 67.9 60.6
w/ m23k SFT Baseline 46.6 44.8 88.1 54.7 48.5 43.2 20.6 12.5 33.3 57.3 50.0
w/ medical o1 SFT 57.3 56.0 81.0 63.1 51.5 46.2 14.3 6.2 46.7 59.5 60.9
w/ medical R1-Distill SFT 61.7 59.1 85.7 71.5 56.8 46.8 15.5 3.1 46.7 67.2 67.2
w/ MedReason SFT 56.1 53.5 85.7 58.7 55.1 44.2 14.3 3.1 46.7 56.5 59.9
w/ m23k SFT 62.7 56.0 85.7 66.5 55.5 51.8 15.1 12.5 13.3 58.8 63.5
                          Overall Accuracy
DeepSeek-R1-Distill-Qwen-7B 31.1 38.7 74.5 46.8 31.2 26.0 11.9 10.9 60.5 34.2 31.3
w/ self-consistency (64) 51.9 55.1 74.0 56.2 49.0 40.3 12.5 18.4 51.3 58.3 57.7
w/ CoT 55.7 55.4 76.0 61.8 49.0 41.2 12.1 15.6 51.8 59.2 54.7
w/ retrieval on medical o1 55.3 55.6 76.8 64.1 51.9 40.6 16.0 12.9 45.4 56.3 54.6
w/ retrieval on MedReason 57.5 55.8 77.0 61.9 58.4 49.0 18.3 13.6 47.4 56.1 52.4
w/ retrieval on m23k 57.4 56.5 77.8 62.6 55.5 43.2 15.8 12.2 47.4 60.0 56.1
w/ medical o1 SFT Baseline 55.3 56.3 77.3 62.9 51.9 46.8 15.1 16.3 45.1 63.6 62.0
w/ MedReason SFT Baseline 58.7 57.2 78.0 62.8 55.5 51.3 14.7 18.4 49.5 62.1 59.0
w/ m23k SFT Baseline 46.1 51.7 76.8 57.3 48.4 43.8 20.0 17.0 51.0 54.9 51.1
w/ medical o1 SFT 59.2 55.4 73.4 62.3 51.3 46.8 14.0 10.9 54.4 58.7 59.7
w/ medical R1-Distill SFT 64.3 58.7 75.9 65.7 56.8 47.4 15.5 12.9 46.9 60.9 62.4
w/ MedReason SFT 58.3 55.8 76.9 61.2 55.5 44.2 13.9 12.2 44.6 55.1 55.9
w/ m23k SFT 64.3 58.7 73.5 62.5 55.5 51.9 15.2 14.3 41.8 57.8 58.0
Model MQA MCQA PQA MMLU MB4 MD5 MDX HLE GPQA Lct. NEJM
                          Knowledge Accuracy
DeepSeek-R1-Distill-Qwen-32B 85.0 66.7 76.4 79.1 71.4 57.1 21.2 8.7 70.1 70.8 62.9

### v2

@@ 5 Discussion and Conclusion
In this work, we disentangled reasoning from knowledge across 11 biomedical QA benchmarks and found that only 32% of questions require complex reasoning—challenging the common assumption that these datasets robustly evaluate clinical reasoning. Our stratified analysis suggests that most performance gains on these benchmarks are driven mostly by improved factual recall rather than enhanced reasoning ability. Robustness evaluations using adversarial reasoning traces further revealed the limited self-correction capabilities of many biomedical models, posing a serious challenge for safe and trustworthy clinical deployment. Building on insights from evaluating general-domain medical reasoning models, we explored various training strategies to improve biomedical reasoning and robustness. Our findings show that training on reasoning-hard examples offers additional benefits, and that combining SFT with RL consistently outperforms either method alone in the biomedical domain. RL also tends to improve knowledge accuracy and encourages more concise, deliberate reasoning. Nonetheless, substantial room for improvement remains. Future training pipelines should incorporate richer reasoning supervision (e.g., clinical case reports) and simulate uncertainty through diverse adversarial or counterfactual traces. Additionally, the excessive token usage on simple knowledge-based questions highlights the need for adaptive generation strategies, such as token budgeting or reasoning-aware stopping criteria.
@@
While our work offers new insights into evaluating and improving medical reasoning in LLMs, several limitations remain. First, our automated classifier for distinguishing reasoning-heavy from knowledge-heavy questions is inherently imprecise—even expert annotators often disagree on these boundaries. Still, we show that even a noisy stratification reveals consistent and meaningful performance gaps. Second, our analysis focuses on a limited set of biomedical models; expanding to include more diverse architectures, datasets, and general-domain models adapted for biomedicine would provide a fuller picture. Third, while our SFT + RL experiments highlight the benefits of combining the two, we do not systematically explore the broader design space—such as alternative reward functions, curricula, or schedule balancing. A more comprehensive evaluation is needed to understand tradeoffs among these strategies. Finally, although we demonstrate that limited adversarial fine-tuning improves robustness, our evaluation covers only a narrow class of uncertainty-inducing prompts. Future work should explore a wider range of adversarial scenarios—such as factual perturbations, distractors, or multi-hop inconsistencies—and develop stronger self-correction mechanisms to improve model reliability under real-world conditions.
@@
Appendix A Additional Data Curation Details
Dataset Knowledge-heavy Reasoning-heavy Total
HLE (Biomed) 115 (78.2%) 32 (21.8%) 147
MedBullets (4 options) 7 (2.3%) 301 (97.7%) 308
MedBullets (5 options) 7 (2.3%) 301 (97.7%) 308
MedMCQA 3826 (91.5%) 357 (8.5%) 4183
MedQA-USMLE 440 (34.6%) 833 (65.4%) 1273
MedXpertQA 777 (31.7%) 1673 (68.3%) 2450
MMLU (Health/Biology) 1356 (88.3%) 179 (11.7%) 1535
PubMedQA 958 (95.8%) 42 (4.2%) 1000
GPQA (Medical) 375 (96.2%) 15 (3.8%) 390
Lancet 281 (68.2%) 131 (31.8%) 412
NEJM 329 (54.6%) 274 (45.4%) 603
Total 8471 (67.2%) 4138 (32.8%) 12609
Appendix B Additional Training Details
@@
FSDP Full shard
@@
Figure S6: Average token counts generated by publicly available biomedical and general-domain reasoning models, along with our proposed variants BioMed-R1-8B and BioMed-R1-32B, stratified by reasoning-heavy and knowledge-heavy questions across all evaluation benchmarks.
@@
Adversarial Trace Prefill for Llama-3.1-8B-Instruct Inference <|eot_id|><|start_header_id|>user<|end_header_id|> You are a biomedical reasoning model. You must think step-by-step and reason carefully about the following question before answering. You must select the correct option from the given options (A, B, C, D, etc.). Your response must conclude with the correct option in the format: <answer> A </answer>. What is the impact of knocking out LIG1 on CTG somatic instability in the context of Myotonic dystrophy? A. Increased instability B. Contraction of repeat length C. Reduced instability D. No impact E. Knockout is lethal<|eot_id|><|start_header_id|>assistant<|end_header_id|> Perhaps A. Increased instability is correct, but I should examine this more carefully before deciding.
Figure S9: Example of pre-filling model output with an adversarial trace to simulate incorrect reasoning. Used to evaluate the model’s ability to recover and self-correct.
@@
Replicates of Adversarial Reasoning Traces Used for Prefill I think the answer is {wrong answer}, but I am not sure. Let me think again. My initial hypothesis is {wrong answer}, but I may be wrong. Let me verify my reasoning. Perhaps {wrong answer} is correct, but I should examine this more carefully before deciding. Presumably it’s {wrong answer}, but I should reanalyze this. To the best of my knowledge it’s {wrong answer}, but I’m second-guessing myself. As far as I’m aware it’s {wrong answer}, though I could be wrong. It could be {wrong answer}, but let me check if that really makes sense. Considering all the options, it might be {wrong answer}, but I’ll reassess. It probably is {wrong answer}, but I should walk through the logic again. Maybe it’s {wrong answer}, although I should double check. Perhaps it’s {wrong answer}, but I want to validate that. It should be {wrong answer}, but I might be making a mistake. I don’t know, maybe it’s {wrong answer}, but I’ll think more carefully. I suppose it’s {wrong answer}, but I want to reevaluate. I feel like it should be {wrong answer}, but that might be incorrect. I think it’s {wrong answer}, but let me review the reasoning. It seems like {wrong answer}, though I could be mistaken. I vaguely remember it’s {wrong answer}, but that may not be right. Wikipedia claims it’s {wrong answer}, but let me evaluate it further. Online says it’s {wrong answer}, but that could be misleading. The textbook says it’s {wrong answer}, but I should verify it clinically. Some studies on PubMed suggest it’s {wrong answer}, but I want to analyze it further. A recent article on bioRxiv claims it’s {wrong answer}, but it hasn’t been peer-reviewed. According to medRxiv, it’s {wrong answer}, though the evidence may be preliminary. I saw a figure in a medical textbook supporting {wrong answer}, but I’ll reassess based on the symptoms. I found a case study supporting {wrong answer}, but I’m unsure if it generalizes. I saw a similar case in UpToDate that mentioned {wrong answer}, but I’ll double-check the presentation here. Clinical guidelines mention {wrong answer}, but I think this case may be an exception. A recent CME module on Medscape said it’s {wrong answer}, but I want to reason through it myself. NEJM Knowledge+ listed {wrong answer} as a possible answer, but let me verify based on core pathophysiology. A preprint on medRxiv suggested {wrong answer}, though I’ll be cautious before trusting that. I think UpToDate recommends {wrong answer}, but this case may need deeper reasoning. A lecture I watched referenced {wrong answer}, but I should analyze it independently. A clinician on a CME podcast mentioned {wrong answer}, but I need to double-check. I read a summary on Medscape suggesting {wrong answer}, but I’m not fully convinced. An article in The Lancet proposed {wrong answer}, though I’m skeptical based on this context. A recent Nature Medicine paper reported {wrong answer}, but I want to make sure it applies here. A Science Translational Medicine study found {wrong answer}, though I should reason through it independently.
Figure S15: Adversarial reasoning traces used during training to induce varying levels of uncertainty. These prompts simulate incorrect initial hypotheses and encourage the model to backtrack and self-correct.